# Benchmarking pig detection and tracking under diverse and challenging conditions

### v1

@@ 1 Introduction
Monitoring individual behavior in pig farming is essential to ensure a high level of animal welfare and the efficient functioning of work processes. Traditionally, the acquisition of behavioral information has been time-consuming and laborious, as it relied on human observation and documentation. In the last decade, this has started to change due to the emergence of powerful machine learning methods that allow to automate this task. Automatic methods for individual identification [1, 2] and individualized action understanding [3, 4] have the potential to serve as digital assistance tools for farmers and researchers to monitor the health status [5, 6], to detect potentially harmful behaviors like tail biting[7] or mounting[8, 9], or to observe normal behavior with the goal of detecting changes as early warning signs [10, 11, 12]. At the core of such methods lies the localization of individual animals across space and time, commonly referred to as object detection and multi-object tracking in the technical literature. Since these tasks often are modular components in more sophisticated behavioral analysis pipelines [13], establishing strong, robust and readily available methods for pig detection and tracking is a crucial prerequisite for further developments. While detection and tracking methods based on fine-grained instance representations such as segmentation masks [14, 15], or keypoints exist [16, 17, 1], most of the technical literature is based on axis-aligned bounding boxes. Similarly, state-of-the-art computer vision methods for individualized action understanding primarily rely on bounding boxes [18, 19, 20, 21]. This predominance can be attributed to the ease of annotation and the availability of increasingly powerful backbones [20, 22] that allow tracking- and action-related information to be extracted directly from images or videos without the need for complex intermediate representations. Although rotated bounding boxes [23, 24] can more accurately capture the extent of objects, they are incompatible with the vast majority of multi-object tracking and action understanding methods. As a result, it is difficult to benefit from future advances in these areas when using this representation. For these reasons, our work focuses on detection and tracking methods based on axis-aligned bounding boxes.
@@
Many studies in recent years have addressed the problem of pig detection [5, 11, 25, 26, 27, 28, 29, 30, 31] and tracking [32, 33, 5, 34, 35, 10, 36, 28, 37, 30, 38, 39, 40, 41, 42, 43], often adapting standard methods to better suit the conditions of pig farming, and achieving promising results. Unfortunately, only a few authors made their detection [11, 25, 5, 30] and tracking datasets [10, 30, 40, 42] publicly available. These datasets are limited in terms diversity, often including only a single barn or pen environment, and have not become well established as benchmark datasets, as they are rarely used for comparisons across different works. Other works state that research data is available upon request. However, we did not receive responses to the inquiries we made, suggesting that access to such data may be limited in practice. Similarly, user-friendly code bases for pig detection and tracking are scarce, as research code is often not made publicly available within the livestock research community. Recent works on pig detection and tracking are beginning to acknowledge this gap by making both their code and datasets publicly available [30, 42].
@@
The lack of publicly accessible code bases and datasets is, in our opinion, problematic for two reasons: (1) Current methods cannot be compared to determine the state of the art. Many studies on pig detection [26, 27, 28, 29, 31] and tracking [32, 33, 5, 34, 36, 28, 37, 38, 39, 41] report evaluation metrics on undisclosed datasets, making direct comparisons between them impossible. In the broader field of computer vision, methods are usually benchmarked on standardized, publicly available datasets [e.g. 44, 45, 18], which facilitates reproducibility and encourages fair performance comparisons. (2) The lack of accessible resources also hampers the development and the availability of robust pig detection and tracking methods which are urgently needed for downstream analysis tasks. Researchers that need to localize animals as part of their research often start by re-inventing the wheel. They annotate large amounts of data and adjust generic code bases for detection and tracking to train and hyperparameter-tune their own models without benefiting from the work of researchers that already tackled this task. In contrast, open data and open source are common practice in the broader field of computer vision, such that methods can be easily fine-tuned, extended or used as modular components in custom analysis pipelines. The widespread availability of open-source code bases and high-quality labeled data has been a major driver of the advancements in computer vision in the last decade, and would greatly benefit precision livestock farming research as well.
@@
We benchmarked the performance of several recent general-purpose detection and tracking models on our datasets. In addition to models that are commonly used in pig farming, such as YOLO variants for object detection or SORT-based methods for multi-object tracking, we also evaluated models that have not yet been employed in this context. Specifically, we included Co-DINO [46], a detection model that achieves state-of-the-art results on the COCO dataset [45], and two recent end-to-end trainable models for multi-object tracking [47, 48] that achieve state-of-the-art results on the DanceTrack dataset [49]. Our comprehensive benchmarking study enables a systematic comparison between methods and highlights strengths and weaknesses of different model classes. For the end-to-end models we also conducted a detailed error analysis, providing guidance for future improvements.
We further evaluated all tracking models on a third-party test video from a previously unseen pen environment [42], achieving a substantial improvement over the results reported by the authors. This highlights the importance of carefully selected, high-quality training data for the development of robust and generalizable models. PigDetect (https://doi.org/10.25625/I6UYE9) and PigTrack (https://doi.org/10.25625/P7VQTP), along with the trained model weights and the source code for training, evaluation and inference (https://github.com/jonaden94/PigBench), are made publicly available to facilitate reproducibility, re-use and further development. In summary, our study has three main contributions:
@@ 2.1 Annotation
The images in our PigDetect dataset come from previously unpublished video footage recorded in the projects with grant numbers 28RZ372060 (Inno Pig), 2817902315 (FriSch), EIP-Agri-Ni-2019-LPTT-03 (Old Breed New House), and 2822ZR015 (TiPP), as well as image data published in prior works [50, 10, 51] that had not yet been annotated with bounding boxes. To obtain images for annotation, 2363 images were randomly sampled from the available image (N = 975) and video sources (N = 1388). For the video sources, frames were extracted as sparsely as possible to ensure diversity. Specifically, we randomly sampled distinct one-hour videos and subsequently extracted one frame from each of these videos. The sampled frames were then annotated with axis-aligned bounding boxes using the labelme tool [52] and form the foundation of our benchmark dataset.
Some challenging detection scenarios cannot be adequately captured with random sampling because they occur too infrequently. To identify images as “challenging”, we trained a pig detector (YOLOX, [53]) with the randomly sampled images and visually inspected the predicted bounding boxes for a large number of images from all sources using the interactive labeling tool pylabel [54]. If the predicted bounding boxes were considered to contain severe errors (no formal criterion), the annotation was corrected. This way, another 568 images were added to the dataset from the image (N = 252) and video sources (N = 316), resulting in a total of 2931 images. Of these images, 1704 were sampled from the previously unpublished video footage (Inno Pig, FriSch, Old Breed New House, TiPP), and 1227 from the other sources [50, 10, 51].
The annotation was performed by one person and independently reviewed by another, with corrections applied where needed. Two of the authors were involved in this procedure (J. H. and C. P.). It should be noted that some images could not be unambiguously annotated (e. g. cases of heavy piling) and were therefore not included in the dataset, which highlights the practical limitations of pig detection.
@@ 2.2 Characteristics
Figure 1: Distributional properties of ground truth annotations and example images of PigDetect. (a) displays the distribution of the number of pigs per image across all images. The bin width is set to one. (b) shows the distribution of a pig’s maximum intersection over union (IoU) with another pig in the same image across all pigs in all images. The bin width is set to 0.025. (c) shows the distribution of a pig’s bounding box size relative to image size across all pigs in all images. The bin width is set to 0.005. For subfigures b and c, large values that occur too rarely to be visible in the histograms were excluded. (d) depicts example images from PigDetect that include challenging conditions such as occlusions and bad visibility. The top row shows images from Inno Pig, Old Breed New House, TiPP, and FriSch (left to right). The bottom row shows images from Alameer et al. [51], Bergamini et al. [10], and two from Psota et al. [50] (left to right).
Table 1: Dataset characteristics of PigDetect compared to publicly available pig object detection datasets. Previous works did not distinguish between random and challenging images, so the field is left blank. Night images and challenging images are not mutually exclusive.
Dataset #Images #Challenging #Night #BBoxes #Pens
Riekert et al. [11, 25] 582 - 156 13 0201302013\,02013 020 18
Alameer et al. [5] 113 379113379113\,379113 379 - 0 735 094735094735\,094735 094 1
Melfsen et al. [30] 614 - 0 6754 1
PigDetect (our) 2931 568 687 42 1774217742\,17742 177 31
train 2431 414 581 33 1973319733\,19733 197 30
@@
Tab. 1 provides dataset characteristics of PigDetect compared to other publicly available pig detection datasets. Previously published works employed only random sampling for dataset curation without specifically selecting challenging images, which we show to be suboptimal (see Sec. 6.2). Furthermore, PigDetect contains more images and annotated bounding boxes than most previously published datasets [11, 25, 30]. Alameer et al. [5] published an action understanding dataset that contains images from pig farming annotated with bounding boxes that are associated with an action label. While this dataset is more extensive than PigDetect, it contains only daytime images from a single pen. Compared to all other datasets, PigDetect includes a larger number of visually diverse pens (Fig. 1d) that were captured during day and night. A comprehensive depiction of the visual diversity of the pen settings of previously published datasets (Fig. 7) and PigDetect (Fig. 8) can be found in A. For a more detailed breakdown of image characteristics for the different sources included in PigDetect, we refer to Tab. 9.
@@
Image resolutions in our dataset vary greatly, with the smallest and largest images having a width and height of $630\times 355$ pixels and $3840\times 2160$ pixels respectively. This is neither a problem for object detection nor for tracking since training and inference pipelines usually include automatic resizing and padding. For night images, most sources used infrared imaging. In total there are 687 night images in the dataset (23 %). The data coming from one barn of Alameer et al. [51] consists of images where only an infrared camera was used during the day. These 153 images appear as grayscale and were thus counted as night images.
@@
The majority of pigs found in the images from the video sources are crossbreed pigs, (German Large White × German Landrace) × (Piétrain or Duroc). Images from the project “Old Breed New House” contain Swabian Hallish, and Bentheim crossbreed pigs. The animals in the images from Alameer et al. [51] are either Duroc × (Landrace × Large White) or Danbred Duroc × (Danbred Yorkshire × Danbred Landrace). Psota et al. [50] and Bergamini et al. [10] did not disclose the breed of their pigs, but they appear to be mostly typical crossbreeds, with a few exceptions in Psota et al. [50] that include pigs with spotted or belted colors, as well as a few pigs that appear to be Poland China pigs. Throughout the dataset, pigs were captured at different stages of development from rearing to fattening.
@@
Images taken from Psota et al. [50] are the most diverse in terms of backgrounds, containing 17 different pens. The images in Alameer et al. [51] are from multiple pens in two different barn locations, and in Bergamini et al. [10] only a single pen background is present. The data from Inno Pig, FriSch, Old Breed New House and TiPP (see Sec. 2.1) contains 11 distinct environments that differ in floor type and pen equipment.
@@
The pens were captured from above, often with multiple cameras at different angles. Most of the images were taken in a practical barn environment. This resulted in various conditions and image characteristics that pose a challenge to object detectors. These include low contrast due to insufficient lighting or overexposure, smudges or dust on the camera lens, motion blur, and partial occlusion of pigs through other pigs or barn interior. One reason for severe occlusions is piling of pigs, which is also included in the dataset to the extent that it could still be annotated. Some images do not fully capture the pen they are monitoring, which results in pigs that are cut off at the edges of the image. Additionally, some images contain parts of other pens. If possible, these pigs were annotated as well.
@@
The distribution of the number of pigs in an image has local maxima around 10 and 23 (Fig. 1a), with the minimum and maximum number of pigs being one and 32. The bounding boxes of a large number of pigs have little to no overlap with those of other pigs, reflecting the fact that pigs are often lying or standing around alone to sleep, rest, eat or drink. However, the dataset also contains a large amount of pigs whose maximum intersection over union (IoU) with the bounding box of another pig is substantially larger than zero (Fig. 1b). Bounding box sizes relative to the image size vary widely from very small bounding boxes that comprise less than 1 % of the image up to larger bounding boxes that comprise 10 % of the image (Fig. 1c). Example images from diverse barn environments that also showcase challenging detection scenarios are depicted in Fig. 1d.

### v2

@@ 1 Introduction
Monitoring individual behavior in pig farming is essential to ensure a high level of animal welfare and the efficient functioning of work processes. Traditionally, the acquisition of behavioral information has been time-consuming and laborious, as it relied on human observation and documentation. In the last decade, this has started to change due to the emergence of powerful machine learning methods that allow to automate this task. Automatic methods for individual identification [t_psota_long-term_2020, wutke2025multistage] and individualized action understanding [tong2022videomae, wu2021towards] have the potential to serve as digital assistance tools for farmers and researchers to monitor the health status [alameer_automated_2020, bhujel_deep-learning-based_2021], to detect potentially harmful behaviors like tail biting[liu_computer_2020] or mounting[yang2021pig, li_mounting_2019], or to observe normal behavior with the goal of detecting changes as early indicators of health and welfare problems [bergamini_extracting_2021, riekert_automatically_2020, hesse2023computer]. At the core of such methods lies the localization of individual animals across space and time, commonly referred to as object detection and multi-object tracking in the technical literature. Since these tasks often are modular components in more sophisticated behavioral analysis pipelines [vogg2025computer], establishing strong, robust and readily available methods for pig detection and tracking is a crucial prerequisite for further developments. While detection and tracking methods based on fine-grained instance representations such as segmentation masks [he2017mask, wang_end--end_2021], or keypoints exist [toshev2014deeppose, mathis2018deeplabcut, t_psota_long-term_2020], most of the technical literature is based on axis-aligned bounding boxes. Similarly, state-of-the-art computer vision methods for individualized action understanding primarily rely on bounding boxes [gu2018ava, gritsenko_end--end_2023-1, ryali2023hiera, wu_memvit_2022]. This predominance can be attributed to the ease of annotation and the availability of increasingly powerful backbones [ryali2023hiera, dosovitskiy_image_2021] that allow tracking- and action-related information to be extracted directly from images or videos without the need for complex intermediate representations. Although rotated bounding boxes [liu2017learning, liu2023s] can more accurately capture the extent of objects, they are incompatible with the vast majority of multi-object tracking and action understanding methods. As a result, it is difficult to benefit from future advances in these areas when using this representation. For these reasons, our work focuses on detection and tracking methods based on axis-aligned bounding boxes.
@@
Many studies in recent years have addressed the problem of pig detection [alameer_automated_2020, riekert_automatically_2020, riekert2021model, hao_novel_2022, ji_automatic_2022, guo_enhanced_2023, mattina_efficient_2023, melfsen2023describing, pu2024tr] and tracking [cowton_automated_2019, zhang2019automatic, alameer_automated_2020, gan_automated_2021, shirke2021tracking, bergamini_extracting_2021, wang_towards_2022, guo_enhanced_2023, parmiggiani2023don, melfsen2023describing, wei2023detection, yang2024long, jaoukaew2024robust, tu2024tracking, yu2025fto, liu2025sdgtrack], often adapting standard methods to better suit the conditions of pig farming, and achieving promising results. Unfortunately, only a few authors made their detection [riekert_automatically_2020, riekert2021model, alameer_automated_2020, melfsen2023describing] and tracking datasets [bergamini_extracting_2021, melfsen2023describing, jaoukaew2024robust, yu2025fto] publicly available. These datasets are limited in terms diversity, often including only a single barn or pen environment, and have not become well established as benchmark datasets, as they are rarely used for comparisons across different works. Other works state that research data is available upon request. However, we did not receive responses to the inquiries we made, suggesting that access to such data may be limited in practice. Similarly, user-friendly code bases for pig detection and tracking are scarce, as research code is often not made publicly available within the livestock research community. Recent works on pig detection and tracking are beginning to acknowledge this gap by making both their code and datasets publicly available [melfsen2023describing, yu2025fto].
@@
The lack of publicly accessible code bases and datasets is, in our opinion, problematic for two reasons: (1) Current methods cannot be compared to determine the state of the art. Many studies on pig detection [hao_novel_2022, ji_automatic_2022, guo_enhanced_2023, mattina_efficient_2023, pu2024tr] and tracking [cowton_automated_2019, zhang2019automatic, alameer_automated_2020, gan_automated_2021, wang_towards_2022, guo_enhanced_2023, parmiggiani2023don, wei2023detection, yang2024long, tu2024tracking] report evaluation metrics on undisclosed datasets, making direct comparisons between them impossible. In the broader field of computer vision, methods are usually benchmarked on standardized, publicly available datasets [e. g. deng2009imagenet, lin2014microsoft, gu2018ava], which facilitates reproducibility and encourages fair performance comparisons. (2) The lack of accessible resources also hampers the development and the availability of robust pig detection and tracking methods which are urgently needed for downstream analysis tasks. Researchers that need to localize animals as part of their research often start by re-inventing the wheel. They annotate large amounts of data and adjust generic code bases for detection and tracking to train and hyperparameter-tune their own models without benefiting from the work of researchers that already tackled this task. In contrast, open data and open source are common practice in the broader field of computer vision, such that methods can be easily fine-tuned, extended or used as modular components in custom analysis pipelines. The widespread availability of open-source code bases and high-quality labeled data has been a major driver of the advancements in computer vision in the last decade, and would greatly benefit precision livestock farming research as well.
@@
We benchmarked the performance of several recent general-purpose detection and tracking models on our datasets. In addition to models that are commonly used in pig farming, such as YOLO variants for object detection or SORT-based methods for multi-object tracking, we also evaluated models that have not yet been employed in this context. Specifically, we included Co-DINO [zong2023detrs], a detection model that achieves state-of-the-art results on the COCO dataset [lin2014microsoft], and two recent end-to-end trainable models for multi-object tracking [zhang_motrv2_2023, gao2025multiple] that achieve state-of-the-art results on the DanceTrack dataset [sun2022dancetrack]. Our comprehensive benchmarking study enables a systematic comparison between methods and highlights strengths and weaknesses of different model classes. For the end-to-end models we also conducted a detailed error analysis, providing guidance for future improvements.
We further evaluated both the detection and tracking models on a third-party dataset from a previously unseen pen environment, where they demonstrated strong performance. In particular, the tracking models substantially outperformed the best results previously reported by yu2025fto. This highlights the importance of carefully selected, high-quality training data for the development of robust and generalizable models. PigDetect (https://doi.org/10.25625/I6UYE9) and PigTrack (https://doi.org/10.25625/P7VQTP), along with the trained model weights and the source code for training, evaluation and inference (https://github.com/jonaden94/PigBench), are made publicly available to facilitate reproducibility, re-use and further development. In summary, our study has three main contributions:
@@ 2.1 Annotation
The images in our PigDetect dataset come from previously unpublished video footage recorded in the projects with grant numbers 28RZ372060 (Inno Pig), 2817902315 (FriSch), EIP-Agri-Ni-2019-LPTT-03 (Old Breed New House), and 2822ZR015 (TiPP), as well as image data published in prior works [psota_multi-pig_2019, bergamini_extracting_2021, alameer_automated_2022] that had not yet been annotated with bounding boxes. By including third-party sources, we aimed to further increase the diversity of the dataset. To obtain images for annotation, 2363 images were randomly sampled from the available image (N = 975) and video sources (N = 1388). For the video sources, frames were extracted as sparsely as possible to ensure diversity. Specifically, we randomly sampled distinct one-hour videos and subsequently extracted one frame from each of these videos. The sampled frames were then annotated with axis-aligned bounding boxes using the labelme tool [wada_labelme_2023] and form the foundation of our benchmark dataset.
Some challenging detection scenarios cannot be adequately captured with random sampling because they occur too infrequently. To identify images as “challenging”, we trained a pig detector (YOLOX, [ge_yolox_2021]) with the randomly sampled images and visually inspected the predicted bounding boxes for a large number of images from all sources using the interactive labeling tool pylabel [heaton_pylabel_2023]. If the predicted bounding boxes were considered to contain severe errors (no formal criterion), the annotation was corrected. The most frequent error cases were missed detections (e. g. due to occlusion or poor visibility), multiple bounding boxes for a single pig (e. g. for a pig split into two separate visible parts through occlusion), and highly inaccurate bounding boxes (e. g. two pigs erroneously detected as one in low-contrast scenarios). By identifying and correcting erroneous cases, another 568 images were added to the dataset from the image (N = 252) and video sources (N = 316), resulting in a total of 2931 images. Of these images, 1704 were sampled from the previously unpublished video footage (Inno Pig, FriSch, Old Breed New House, TiPP), and 1227 from the other sources [psota_multi-pig_2019, bergamini_extracting_2021, alameer_automated_2022].
The annotation was performed by one person and independently reviewed by another, with corrections applied where needed. Two of the authors were involved in this procedure (J. H. and C. P.). It should be noted that some images could not be unambiguously annotated (e. g., cases of heavy piling) and were therefore not included in the dataset, which highlights the practical limitations of pig detection.
@@ 2.2 Characteristics
Figure 1: Distributional properties of ground truth annotations and example images of PigDetect. (a) displays the distribution of the number of pigs per image across all images. The bin width is set to one. (b) shows the distribution of a pig’s maximum intersection over union (IoU) with another pig in the same image across all pigs in all images. The bin width is set to 0.025. (c) shows the distribution of a pig’s bounding box size relative to image size across all pigs in all images. The bin width is set to 0.005. For subfigures b and c, large values that occur too rarely to be visible in the histograms were excluded. (d) depicts example images from PigDetect that include challenging conditions such as occlusions and bad visibility. The top row shows images from Inno Pig, Old Breed New House, TiPP, and FriSch (left to right). The bottom row shows images from alameer_automated_2022, bergamini_extracting_2021, and two from psota_multi-pig_2019 (left to right).
Table 1: Dataset characteristics of PigDetect compared to publicly available pig object detection datasets. Previous works did not distinguish between random and challenging images, so the field is left blank. The number of images that are both challenging and captured during night is 97 (train), 34 (val) and 27 (test).
Dataset Images Challenging Night BBoxes Pens
riekert_automatically_2020, riekert2021model 582 - 156 13 02013\,020 18
alameer_automated_2020 113 379113\,379 - 0 735 094735\,094 1
melfsen2023describing 614 - 0 6754 1
PigDetect (our) 2931 568 687 42 17742\,177 31
train 2431 414 581 33 19733\,197 30
@@
Tab. 1 provides dataset characteristics of PigDetect compared to other publicly available pig detection datasets. Previously published works employed only random sampling for dataset curation without specifically selecting challenging images, which we show to be suboptimal (see Sec. 6.3). Furthermore, PigDetect contains more images and annotated bounding boxes than most previously published datasets [riekert_automatically_2020, riekert2021model, melfsen2023describing]. alameer_automated_2020 published an action understanding dataset that contains images from conventional pig farming annotated with bounding boxes that are associated with an action label. While this dataset is more extensive than PigDetect, it contains only daytime images from a single pen. Compared to all other datasets, PigDetect includes a larger number of visually diverse pens from conventional indoor barn environments (Fig. 1d) that were captured during day and night. A comprehensive depiction of the visual diversity of the pen settings of previously published datasets (Fig. 7) and PigDetect (Fig. 8) can be found in A. For a more detailed breakdown of image characteristics for the different sources included in PigDetect, we refer to Tab. 10.
@@
Image resolutions in our dataset vary greatly, with the smallest and largest images having a width and height of $630\times 355$ pixels and $3840\times 2160$ pixels respectively. This is neither a problem for object detection nor for tracking since training and inference pipelines usually include automatic resizing and padding. For night images, most sources used infrared imaging. In total, there are 687 night images in the dataset (23 %). The data coming from one barn of alameer_automated_2022 consists of images where only an infrared camera was used during the day. These 153 images appear as grayscale and were thus counted as night images.
@@
The majority of pigs found in the images from the video sources are crossbreed pigs, (German Large White × German Landrace) × (Piétrain or Duroc). Images from the project “Old Breed New House” contain Swabian Hallish, and Bentheim crossbreed pigs. The animals in the images from alameer_automated_2022 are either Duroc × (Landrace × Large White) or Danbred Duroc × (Danbred Yorkshire × Danbred Landrace). psota_multi-pig_2019 and bergamini_extracting_2021 did not disclose the breed of their pigs, but they appear to be mostly typical crossbreeds, with a few exceptions in psota_multi-pig_2019 that include pigs with spotted or belted colors, as well as a few pigs that appear to be Poland China pigs. Throughout the dataset, pigs were captured at different stages of development from rearing to fattening.
@@
Images taken from psota_multi-pig_2019 are the most diverse in terms of backgrounds, containing 17 different pens. The images in alameer_automated_2022 are from multiple pens in two different barn locations, and in bergamini_extracting_2021 only a single pen background is present. The data from Inno Pig, FriSch, Old Breed New House and TiPP (see Sec. 2.1) contains 11 distinct environments that differ in floor type and pen equipment.
@@
The pens were captured from above, often with multiple cameras at different angles. Most of the images were taken in a practical barn environment. This resulted in various conditions and image characteristics that pose a challenge to object detectors. These include low contrast due to insufficient lighting or overexposure, smudges or dust on the camera lens, motion blur, and partial occlusion of pigs through other pigs or barn interior. One reason for severe occlusions is piling of pigs, which is also included in the dataset to the extent that it could still be annotated. Some images do not fully capture the pen they are monitoring or contain parts of other pens, which results in pigs that are only partially visible at the edges of the image. If these pigs could still be clearly identified, they were annotated as well.
@@
The distribution of the number of pigs in an image has local maxima around 10 and 23 (Fig. 1a). The minimum number of pigs per video is one, and the maximum is 32. To quantify the extent of overlap between pigs, we computed for each bounding box the maximum intersection-over-union (IoU) with any other bounding box (Fig. 1b). The resulting distribution shows that the bounding boxes of a large number of pigs have little to no overlap with those of other pigs, reflecting the fact that pigs are often lying or standing around alone to sleep, rest, eat or drink. However, it is also evident that most pigs in the dataset have bounding boxes that overlap with those of others. Bounding box sizes relative to the image size vary widely from very small bounding boxes that comprise less than 1 % of the image up to larger bounding boxes that comprise 10 % of the image (Fig. 1c). Example images from diverse barn environments that also showcase challenging detection scenarios are depicted in Fig. 1d.